# Amuro & Char: Analyzing the Relationship between Pre-Training and Fine-Tuning of Large Language Models

### v2

@@ 3.3 Evaluation
Figure 3: Example of few-shot performance on different pre-training steps of the models that benefited (3(a)) and did not benefit from fine-tuning (3(b)). The solid blue line represents the fine-tuned checkpoint, and the dashed orange line represents the base checkpoint. The results of all datasets can be found in Figure 9 and Figure 8.
@@
Figure 4: Amount of performance increase brought by fine-tuning between tasks that model can solve in pre-training (mandarin orange) and tasks that the model could not solve until fine-tuning (sage green). The exact number of mean increase is shown in Appendix G.
4 How does the model change across pre-training?
We begin our evaluation by considering how additional pre-training changes the BASE model. Typically, researchers track the value of the training or held-out loss during training. However, performance improvements on downstream tasks do not always follow the same trend with the loss curves Groeneveld et al. (2024).
We evaluate the pre-trained checkpoints with few-shot examples, as models without alignment tend to do poorly in a zero-shot context. Four shots are randomly sampled from the datasets, which are selected based on the highest performance shot amount reported in Yang et al., 2024. The model’s performance at each pre-training step is reported in Figure 2.
Broadly speaking, our results suggest that all datasets fall into one of two groups. For the first group of datasets (Figure 2(a)), although the model shows clear improvement during the early stages of pre-training, performance levels off fairly early on and remains consistent. The dramatic improvements in the early stages of pre-training may result from larger steps in early optimization. We find improvements stop increasing past step 342,000. The second group (Figure 2(a)) shows tasks that are never learned during pre-training. Performance remains constant throughout the whole pre-training process. These datasets include MNLI, XSum, and BoolQ, and we found no difference between zero-shot and few-shot evaluations. A natural hypothesis for this finding is potential data contamination in the pre-training data. However, the evaluation datasets are selected based on the popularity of the task and the content of pre-training data. All datasets that experience improvement do not exist in the model’s pre-training data Soldaini et al. (2024), while the more likely leaked datasets (MNLI, XSUM) never gain an improvement during the pre-trining process.
Overall, these results reveal an interesting dichotomy. Some tasks can be learned during pre-training, while others are not. Next, we explore what exactly the model is learning regarding this second group of datasets during pre-training by exploring the fine-tuned models.
@@ 4 How does the model change across pre-training?
5 Does more pre-training improve fine-tuning?
@@
Groeneveld et al., 2024 compares OLMo’s performance on several tasks before and after fine-tuning the final checkpoint and finds that fine-tuning enables the model to do well on tasks for which the unaligned model does poorly. We observe (§4) that while some datasets improved during pre-training, there is a group of datasets for which a pre-trained model does poorly. Does the model learn anything that helps solve these tasks, and is fine-tuning required to do well on them? Alternatively, does the model learn useful information for these tasks but cannot express it without fine-tuning? In this section, we further explore this dataset dichotomy by examining fine-tuned checkpoints for each of the datasets.
@@
Our results appear in Figure 3 and Figure 4. First, we consider those datasets where the pre-trained models do well (Figure 2(a)). These datasets do not improve with fine-tuning, suggesting whatever is learned during fine-tuning, which we discuss below, the model already gains the knowledge during pre-training. This effect is observed at all checkpoints; fine-tuning simply does not help.
However, a different story is observed for datasets that are not learned during pre-training. For these, fine-tuning yields significant improvements at every model checkpoint, with Figure 4 showing the magnitude of improvement on these datasets compared to no improvement to the datasets already learned during pre-training. Moreover, earlier checkpoints obtain more substantial gains from fine-tuning than later checkpoints. The benefit of fine-tuning continues to increase until a certain threshold in pre-training steps is reached (approximately 424,000).
@@
Figure 3 shows representative plots comparing the performance of a pre-trained versus fine-tuned model at different checkpoints for two datasets (full list in Appendix E). For Hellaswag (learned during pre-training), fine-tuning does not benefit the model, even during early checkpoints when the model performs poorly on the task. Nevertheless, for MNLI (not learned during pre-training), fine-tuning dramatically improves the model. Interestingly, later checkpoints achieve better results after fine-tuning, even when the performance of the pre-trained model is unchanged. This suggests that the model is, in fact, learning important information during pre-training, but it cannot express that information without fine-tuning.
@@
Our findings suggest that early stopping in pre-training will not be detrimental to downstream fine-tuning performance, and the benefits of fine-tuning an LLM could exceed the benefits of continued pretraining, which sheds light on the potential of cost-effective training paradigm with less pre-training. However, it is difficult to directly identify such a stopping criteria without fine-tuning intermediate checkpoints; the improvement trend is invisible before fine-tuning the checkpoints. Future work may reveal other signals of pre-training behavior that correlate with downstream task performance after fine-tuning. Overall, when resource-intensive pre-trained LLMs are not available, fine-tuning models on models with less pre-training may be a reasonable practical choice for obtaining a high-quality model.
6 Supervised Fine-Tuning: What does the model learn and forget?
What exactly is the model learning during fine-tuning such that it shows abilities in pre-trained models for some tasks but provides no benefit for other tasks? We analyze the supervised fine-tuning process to understand what is learned and what is forgotten. Specifically, we explore three dimensions: task format, task transfer, and domain knowledge.
((a)) MNLI matched
((b)) SocialIQa
6.1 Task Format
Sclar et al., 2023 show that LLMs are extremely sensitive to prompt perturbation in few-shot settings. More broadly, extensive work on prompt engineering reveals the sensitivity of models to task format. We hypothesize that fine-tuning fits the model to a specific task format, resulting in higher performance when the evaluation set matches this format. To test this hypothesis, we vary the task format to either match the training format, use a different format, or rely on instructions. We carefully construct three different prompt formats for the following settings. 1) Default is the same format used for training, where we expect the model to benefit from learning the task format; 2) In contrast, IO format reflects a common way of performing supervised fine-tuning by incorporating only unprocessed input and output; 3) Instruct uses a human-readable instruction template to format the input. Table 4 shows an example of each format. Checkpoint performance before and after fine-tuning is shown in Figure 5.
In the early pre-training steps, aligning the task format with fine-tuning data seems to play a crucial role. The model does not yet have enough information to overcome the differences between the training and test formats. However, when fine-tuned on later pre-training checkpoints, the model gradually becomes more flexible with different task formats, suggesting that model sensitivity to prompt formatting observed may be resolvable with more pre-training and a fine-tuning stage. In this view, fine-tuning teaches the model how to format a response for the task.
6.2 Task Transfer
Numerous studies examine model forgetting, where further model training causes improvements on some tasks but degradation on others Mehta et al. (2023). We evaluate model forgetfulness by examining whether the model does worse on some tasks after fine-tuning for other tasks. Specifically, we divide our tasks into two types: classification and generation. We notate the training datasets as $D_{T}$ and the evaluation datasets as $D_{E}$. We represent the performance of a pre-trained model (BASE) on checkpoint $i$ as ${\text{Perf}}_{BASE}^{i}(d)$ where an evaluation dataset $d\in D_{E}$, and the performance of the i-th checkpoint fine-tuned on dataset $t\in D_{t}$ be $\text{Perf}_{t}^{i}(d)$. To normalize the effect caused by uneven performance across different datasets, we compute the mean ratio of change (MRC) in performance for each checkpoint as follows.
Models fine-tuned on classification tasks and evaluated on generation tasks decrease on average 61.4% compared to models that are never fine-tuned. In contrast, models fine-tuned on generation tasks can still perform the same as the BASE model on classification tasks, with a 0.3% MRC, which is not statistically significantly different from a 0% change. Our findings on all pre-training checkpoints align with the findings of Yang et al. (2024) on the final checkpoint of LLAMA-7B.
Regardless of the pre-training stage, a model can maintain classification abilities when trained for generation, but it loses its generation abilities when trained for classification. This is perhaps not surprising given that classification tasks can be seen as a subset of generation, while the reverse is not true. The model follows a simplicity bias and thus is more likely to memorize simple classification tasks than generation tasks with an exponentially larger search space. Additionally, since we evaluate the classification tasks based on the output logits and the base model performs randomly on the classification tasks, it is much easier for the models to maintain the same performance as the BASE models. Fine-tuning can cause a model to lose abilities when the desired fine-tuning behavior does not support those abilities.
((a)) Paws $\rightarrow$ QQP
((b)) MNLI $\rightarrow$ GPT3NLI
Figure 7: Ratio of out-of-domain performance change for each task, averaged across checkpoints
6.3 Domain Knowledge
Finally, we explore how a model’s generalization ability is affected by fine-tuning by inspecting whether the model forgets the domain knowledge it had before fine-tuning due to learning other abilities. An example of OOD model performance is shown in Figure 6, and the mean change ratio by datasets is presented in Figure 7.
The model does not benefit equally from the in-domain fine-tuning: all NLI datasets experience a boost when fine-tuning on MNLI, while fine-tuning on Paws is detrimental to other paraphrase detection datasets. This implies that both forgetting and learning are happening: the model learns to perform the task with in-domain knowledge, but it may, in turn, forget information more distant from what is learned in fine-tuning. Questions remain, however, about whether there are different stages of learning and forgetting during fine-tuning and whether the model picks up different tasks in various stages, which requires further study of fine-tuning dynamics.
Overall, across these three lenses, we find that fine-tuning, although teaches a model how to perform a task, can sacrifice generalization abilities if such ability is not needed for the fine-tuned task. For some datasets learned with pre-training alone, the model can easily understand the task format, and the nature of the task is probably supported by the pre-training objective. For tasks that can only be learned with subsequent fine-tuning, the model may require additional examples to adapt to different task formats, or the task itself may be inconsistent with the pre-training objective.
@@ 7 Discussion
Our study uses fine-tuning of pre-training model checkpoints to understand the dynamics of pre-training and fine-tuning on model performance. While our insights suggest directions for future work, we note important limitations inherent in our experiments. This study considered a single, relatively small LLM on less than a dozen datasets, and still consumed thousands of hours of GPU training time at significant expense. Future work needs to confront these issues on larger models and more datasets. We believe our experiments can focus future work on specific experiments with larger models.
Some datasets can be learned without fine-tuning. We discover a dichotomy between datasets. Some are learned during model pre-training, while others show no improvements during pre-training. Furthermore, the datasets learned during pre-training do not benefit from fine-tuning. This observation, combined with our study about what is learned during fine-tuning (Section 6) suggests that some tasks are presented in a manner that aligns with what the model sees during pre-training, and thus fine-tuning provides no additional information. While we could identify what about the tasks placed them in the learned or not learnable during pre-training group, it may be possible to format tasks in a manner that better aligns with pre-training and makes them learnable.
Pre-training models can improve in undetectable ways without fine-tuning. Some datasets are not learnable during pre-training but benefit significantly from fine-tuning (§4). However, these datasets still benefited from additional pre-training, even though those benefits were not revealed without fine-tuning (§5). Clearly, the model is learning important information about the task, even though it cannot express that information. The identification of a measure available during pre-training that correlated with post-fine-tuning task performance could be used to guide pre-training and produce models that did better post-fine-tuning. Perhaps there is a way in which information about these tasks can be included in pre-training, allowing the model to better utilize the massive amount of pre-training data. For example, early stopping during pre-training could lead to better utilization of limited training resources if we knew when to stop.
@@
Fine-tuning teaches task format but leads to forgetting unused abilities. Our results show that fine-tuning guides the model to understand the format and complete a given task. As this information diminishes, the model’s overall ability improves. However, fine-tuning comes at the expense of other model abilities, such as the capability of performing on tasks or domains that are unrelated to the fine-tuning task. This insight can be helpful in our understanding of the multitask abilities of LLMs, where certain tasks can introduce conflicts during multi-task training Mueller et al. (2022).
@@ 8 Conclusion
Our experiments explore the relationship between fine-tuning and pre-training LLMs. Our findings span from the latent benefits of pretraining to model learning and forgetting during fine-tuning. Our results show that the model can rapidly pick up the datasets that it could not solve during fine-tuning with only a small amount of supervision. In the meantime, we identify the aspects that LLM learns and forgets during supervised fine-tuning: task format, task solution, and domain knowledge. Overall, our results demonstrate the value of analyzing language model training dynamics, and we would like to call for the release of pre-training checkpoints to aid future studies.
@@
We discuss the weaknesses and limitations in the following section.
Computing Resource
Due to computational constraints, we can only conduct experiments on a 1B model and a limited amount of datasets. The amount of GPU hours spent for each experiment in this study is listed in Table 3.
Availbility of Pre-training Checkpoints
This study would benefit significantly from including a broader spectrum of models, but the public pre-training checkpoint releases are limited. Open-source LLMs with intermediate checkpoint release include OLMo Groeneveld et al. (2024), TinyLLAMA, RedPajama-Incite, OpenLM, and Pythia. After a series of preliminary experiments, we select these models’ best-performing and robust families.
Scaling Law
Recent research shows that the model may resemble emergent capability Wei et al. (2022) when scaled to a certain size. Comparatively, Hassid et al., 2024 find that smaller model is capable of outperforming its larger variant when the computing resources is controlled. To avoid potential confounding factors caused by quantization, our experiments are only conducted on the one-billion model, which may, therefore, conceal the emergent capability brought by larger models while at least giving insights about the potential of small models.
Analysis Protocol
Wu et al., 2023 show that the evaluation result may be affected by samples that have been memorized by the model during training instead of revealing the reasoning capability. The only analysis protocol used is the downstream performance of a trained model. More investigation should be done into model internals during pre-training dynamics and how they relate to the effects of fine-tuning.
Training Paradigm
Although multiple tuning strategies exist, to create a fair comparison environment where checkpoints received the same amount of training, models are fine-tuned with a fixed amount of epochs in this work. On different pre-training stages, the model may converge at a different speed. Further study can be done to study the effect of pre-training on different fine-tuning methods or fine-tuning dynamics in different pre-training stages. We only explored the scenario of full-parameter fine-tuning. Whether parameter-efficient fine-tuning or human preference tuning will lead to a different conclusion also remains an open question.
Randomness
In this study, we only assess uncertainty with Bootstrap during evaluation. However, uncertainty may emerge during training, which poses optimizer initialization and data ordering. Due to the computational constraints, we cannot reduce the randomness factor on this angle.

### v5

@@ 4.2 Domain Knowledge
4.3 Task Transfer
Model forgetting occurs when model training on new tasks improves those tasks at the expense of previously trained tasks (Luo et al., 2023; Mehta et al., 2023; Li and Lee, 2024). To understand whether the model will forget a previously known task solution when fine-tuned on a different one, we evaluate model forgetfulness by examining whether the model does worse on some tasks after fine-tuning for other tasks. Specifically, we divide our tasks into two types: classification and generation.
We notate the training datasets as $D_{T}$ and the evaluation datasets as $D_{E}$. We represent the performance of a pre-trained model (BASE) on checkpoint $i$ as ${\text{Perf}}_{BASE}^{i}(d)$ for an evaluation dataset $d\in D_{E}$, and the performance of the i-th checkpoint fine-tuned on dataset $t\in D_{T}$ be $\text{Perf}_{t}^{i}(d)$. To normalize the effect caused by uneven performance across different datasets, we compute the mean ratio of change (MRC) in performance for each checkpoint as follows.
Models fine-tuned on classification tasks and evaluated on generation tasks decrease on average 61.4% compared to models that are never fine-tuned. In contrast, models fine-tuned on generation tasks can still perform the same as the BASE model on classification tasks, with a 0.3% MRC, which is not statistically significantly different from a 0% change. Our findings on all pre-training checkpoints align with the findings of Yang et al. (2024) on the final checkpoint of LLAMA-7B and our experiments on the final checkpoint of Llama3-8B (Appendix G).
Regardless of the pre-training stage, a model maintains classification abilities when trained for generation but loses generation abilities when trained for classification. This is not surprising given that classification tasks can be seen as a subset of generation, while the reverse is not true. The model follows a simplicity bias (Shah et al., 2020) and thus is more likely to memorize simple classification tasks than generation tasks with an exponentially larger search space. Additionally, since we evaluate the classification tasks based on the output logits and the base model performs randomly on the classification tasks, it is much easier for the models to maintain the same performance as the BASE models. Regardless of the stage of pre-training, fine-tuning can cause a model to lose abilities when the desired fine-tuning behavior does not support those abilities.
Across these three experimental settings, we find that fine-tuning teaches a model how to perform a task without hurting the model’s instruction-following ability, but can sacrifice generalization across domains and tasks.
5 How does the model change across pre-training?
Section 4.1 reveals that the effect brought by fine-tuning could be different depending on the amount of pre-training, but how exactly does pre-training affect downstream fine-tuning results? We begin by considering how additional pre-training changes the BASE model. Typically, researchers track the value of the training or held-out loss during training. However, performance improvements on downstream tasks do not always follow the same trend with the loss curves (Groeneveld et al., 2024).
Instead, we evaluate the pre-trained checkpoints with few-shot examples, as models without alignment tend to do poorly in a zero-shot context. Four shots are randomly sampled from the datasets, which are selected based on the highest performance shot amount reported in Yang et al. (2024). The model’s performance at each pre-training step is reported in Figure 6.
Broadly speaking, our results suggest that all datasets fall into one of two groups. For the first group of datasets (Figure 6(a)), although the model shows clear improvement during the early stages of pre-training, performance levels off fairly early on and remains consistent. The dramatic improvements in the early stages of pre-training may result from larger steps in early optimization. We find improvements stop increasing past step 342,000. The second group (Figure 6(a)) shows tasks that are never learned during pre-training. Performance remains constant throughout the whole pre-training process even when we vary shot sizes. These datasets include MNLI, XSum, and BoolQ. A natural hypothesis for this finding is potential data contamination in the pre-training data. However, the evaluation datasets are selected based on the popularity of the task and the content of pre-training data. All datasets that experience improvement do not exist in the model’s pre-training data (Soldaini et al., 2024), while the more likely leaked datasets (MNLI, XSUM) never gain an improvement during the pre-training process.
Overall, these results reveal an interesting dichotomy. Some tasks can be learned during pre-training, while others cannot. Next, we explore what exactly the model is learning regarding this second group of datasets during pre-training by exploring the fine-tuned models.
@@ 5 How does the model change across pre-training?
Figure 7: Example of few-shot performance on different pre-training steps of the models that benefited (7(a)) and did not benefit from fine-tuning (7(b)). The solid blue line represents the fine-tuned checkpoint, and the dashed orange line represents the base checkpoint. The results of all datasets can be found in Figure 10 and Figure 9.
@@
Figure 8: Amount of performance increase brought by fine-tuning between tasks that model can solve in pre-training (mandarin orange) and tasks that the model could not solve until fine-tuning (sage green). The exact number of mean increase is shown in Appendix J.
@@
6 Does more pre-training yield better fine-tuning results?
@@
Groeneveld et al. (2024) compares OLMo’s performance on several tasks before and after fine-tuning the final checkpoint and finds that fine-tuning enables the model to do well on tasks for which the unaligned model does poorly. We observe (§5) that while some datasets improved during pre-training, there is a group of datasets for which a pre-trained model does poorly. Does the model learn useful information for these tasks but cannot express it without fine-tuning? In this section, we further explore this dataset dichotomy by examining fine-tuned checkpoints for each of the datasets.
@@
Our results appear in Figure 7 and Figure 8. First, we consider those datasets where the pre-trained models do well (Figure 6(a)). These datasets do not improve with fine-tuning, suggesting that whatever is learned during fine-tuning, which we discuss below, the model already gains the knowledge during pre-training. This effect is observed at all checkpoints; fine-tuning simply does not help.
@@
However, a different story is observed for datasets that are not learned during pre-training. For these, fine-tuning yields significant improvements at every model checkpoint, with Figure 8 showing the magnitude of improvement on these datasets compared to no improvement to the datasets already learned during pre-training. Moreover, earlier checkpoints obtain more substantial gains from fine-tuning than later checkpoints. The benefit of fine-tuning continues to increase until a certain threshold in pre-training steps is reached (approximately 424,000). Figure 7 shows representative plots comparing the performance of a pre-trained versus fine-tuned model at different checkpoints for two datasets (full list in Appendix F). For Hellaswag (learned during pre-training), fine-tuning does not benefit the model, even during early checkpoints when the model performs poorly on the task. Nevertheless, for MNLI (not learned during pre-training), fine-tuning dramatically improves the model. Interestingly, later checkpoints achieve better results after fine-tuning, even when the performance of the pre-trained model is unchanged. This suggests that the model is, in fact, improving during pre-training, but it cannot express that improvement without fine-tuning.
@@
Our findings suggest that early stopping in pre-training will not be detrimental to downstream fine-tuning performance. When the budget is limited, the benefits of fine-tuning an LLM could exceed the benefits of continued pretraining, which sheds light on the potential of a cost-effective training paradigm with less pre-training. However, directly identifying such stopping criteria without fine-tuning intermediate checkpoints is challenging. We only empirically observed that the point where more pre-training lead to diminishing return on downstream fine-tuning results approximately align with the turning point of few-shot performance in Section 5. Without such a hypothesis, the improvement trend is invisible before fine-tuning the checkpoints. Overall, when resource-intensive pre-trained LLMs are not available, fine-tuning models on checkpoints with less pre-training may be a reasonable practical choice for obtaining a high-quality model.
@@ 7 Discussion
Our study fine-tunes model pre-training checkpoints to understand the dynamics of pre-training and fine-tuning on model performance.
@@
Fine-tuning teaches additional task format but leads to forgetting unused abilities. Our results show that fine-tuning guides the model to understand the format and complete a given task. As this information diminishes, the model’s overall ability improves. Additionally, more pre-training will lead to a model that reacts better to instruction-style prompts, and the ability to interpret such instruction will not be lost when the model is fine-tuned in a different format. However, fine-tuning comes at the expense of other model abilities, such as the capability of solving tasks or domains that are unrelated or weakly related to the fine-tuning task. This insight can be helpful in our understanding of the multitask abilities of LLMs, where certain tasks can introduce conflicts during multi-task training (Mueller et al., 2022).
Some datasets can be learned without fine-tuning. We discover a dichotomy between datasets. Some are learned during model pre-training, while others show no improvements during pre-training. Furthermore, the datasets learned during pre-training do not benefit from fine-tuning. This observation, combined with our study about what is learned during fine-tuning (§4) suggests that some tasks are presented in a manner that aligns with what the model sees during pre-training, and thus fine-tuning provides no additional information. It may be possible to modify tasks to better align with pre-training and thus make them learnable.
Pre-training can improve models in unseen ways. Some datasets are not learned during pre-training but benefit significantly from fine-tuning (§5). However, these datasets still benefit from additional pre-training, even though those benefits are not revealed without fine-tuning (§6). The model learns important information to solve the task, even though it cannot express that information without fine-tuning. We empirically observe that the point where more pre-training lead to diminishing return on downstream fine-tuning results approximately align with the turning point of few-shot performance in Section 5. Future work may identify ways to verify the turning point and detect these improvements during pre-training, which can better guide pre-training choices to produce models that perform better post-fine-tuning. Perhaps there is a way in which information about these tasks can be included in pre-training, allowing the model to better utilize the massive amount of pre-training data. For example, early stopping during pre-training could lead to better utilization of limited training resources if we know when to stop.
8 Related Work
Recent studies identify phase transition of model training (Olsson et al., 2022; Wei et al., 2022), where new capabilities or behaviors suddenly emerge when certain thresholds of model complexity are reached. The aspects of complexity often include model size, amount of training by FLOPs or tokens, and model architecture. Several prior works studied the training dynamics of language models by analyzing the internals of train-from-scratch models (Tirumala et al., 2022; Chen et al., 2023; Tian et al., 2023; Chen et al., 2024; Chang et al., 2024). The results of these works suggest that the behaviors that are often overlooked after training could be valuable signals for model analysis. In addition to train-from-scratch models, Ren and Sutherland (2024) studied the fine-tuning dynamics of language models. This work focuses on the effect of pre-training dynamics on downstream fine-tuning results by fine-tuning intermediate pre-training checkpoints on various tasks. Due to the scarcity of publically accessible intermediate pre-training checkpoints, the effect of fine-tuning at different pre-training stages is largely unexplored. Concurrent work (Snell et al., 2024) also fine-tunes intermediate pre-training checkpoints and finds that supervised fine-tuning results can be used as a signal to predict when emergence occurs, while our findings point out a dichotomy of model behavior on different datasets, with the potential for data-efficient and budget-friendly training by understanding the stages of model training.
@@ 9 Conclusion
We explore the relationship between fine-tuning and pre-training LLMs through fine-tuning multiple pre-training checkpoints of large language models. Our results on 18 datasets and two models provide insights into LLM training. We identify the aspects that LLM learns and forgets during supervised fine-tuning; By analyzing pre-training history, we find that pre-training improves the model in a latent way that is only observable after fine-tuning. The model may excel at some tasks without fine-tuning. However, the model can rapidly learn datasets that it does not demonstrate capabilities during pre-training with a small amount of supervision. Overall, our study highlights the value of analyzing language model training dynamics. We encourage model developers to release pre-training checkpoints to facilitate research on LLM training.
@@
While our insights suggest directions for future work, we note important limitations inherent in our experiments. We discuss the weaknesses and limitations in the following section.
Computing Resource.
Due to computational constraints, we can only conduct checkpointing experiments on a 1B model. We supply the final checkpoint of an 8B model to verify the findings that are shared across checkpoints. The amount of GPU hours spent for each experiment in this study is listed in Table 4.
Model Size and Variant.
For the analysis with intermediate checkpoints, this study considered a single, relatively small LLM, which may, therefore, conceal the emergent capability brought by larger models (Wei et al., 2022). To combat this, we include the final checkpoint of an 8B model from a different model family. Future work needs to confront these issues on larger models and more datasets.
Availbility of Pre-training Checkpoints.
Although Choshen et al. (2024) points out that the behavior of a model can often be predicted with a model with the same architecture but a different family. This study would benefit significantly from including a broader spectrum of models, but the public pre-training checkpoint releases are limited. We list open-source LLMs with intermediate checkpoint release in Appendix A. After a series of preliminary experiments, we select available models’ best-performing and robust families.
Analysis Protocol.
Wu et al. (2023) show that the evaluation result may be affected by samples that have been memorized by the model during training instead of revealing the reasoning capability. The only analysis protocol used in this work is the downstream performance of a trained model. More investigation should be done into model internals during pre-training dynamics and how they relate to the effects of fine-tuning.
Training Paradigm.
Although multiple tuning strategies exist, to create a fair comparison environment where checkpoints receive the same amount of training, models are fine-tuned with a fixed amount of epochs in this work. On different pre-training stages, the model may converge at a different speed. Further study can be done to study the effect of pre-training on different fine-tuning methods or fine-tuning dynamics in different pre-training stages. We only explored the scenario of full-parameter fine-tuning. Whether parameter-efficient fine-tuning or human preference tuning will lead to a different conclusion also remains an open question.
Randomness.
In this study, we only assess uncertainty with Bootstrap during evaluation. However, uncertainty may emerge during training, which poses optimizer initialization and data ordering, the study of which requires an extensive amount of computing resources.